write the expression for each quantity
1. the value in cents of 5 quarters
2. the value in cents of q quarters
3.the number of months in 7 years
4. the number of monthes in y years
5. the number of gallons in 21 quarts
6.the number of gallons in q quarts ?
i dont get it ? plz help

Answer by scott8148(5895)

you know what these things are; you're just not used to the "formal" mathematical relationships

1) c=5(25)
2) c=q(25)
3) m=7(12)
4) m=y(12)
5) g=21/4
6) g=q/4